Stoichiometric Reactor Problem

Consider the following gas phase reactions in a reactor:

\begin{aligned}Ethylene + W ater & → & Ethanol \\ P ropylene + W ater & → & Isopropyl Alcohol \end{aligned}

Methane is an inert compound that is present in the feed to the reactor. The reactor and the feed are at temperature of 590 K and a pressure of 69 bar. The conversion of ethylene to ethanol is 7% and the conversion of propylene to isopropyl alcohol is 0.7%.

1. Write down the mass balance for each species in this reactor.

2. If the feed has the following composition:

Compound  Feed rate (mol/s)
Methane  200
Ethylene  1289
Propylene  268.6
Ethanol  0.56
Diethyl Ether  0.0
Isopropyl Alcohol 0
Water  773.4

what is the composition of the stream exiting the reactor?

3. Suppose the following additional reaction occurs in series:

2Ethanol → DiethylEther + W ater

with a conversion of 5%. What is the composition of the stream exiting the reactor?

4. Verify your results in CHEMCAD.